# ICYM2I: The illusion of multimodal informativeness under missingness

### v2

@@ 4.1 A motivating example
4.2 ICYM2I-learn: Estimating performances for informativeness under missingness
Training. Under the MAR assumption, i.e., the missingness is fully explained by observed covariates $C$; that is, the probability of a data point being missing depends only on $C$, we propose to train the model with a weighted loss using samples from $\Omega_{\text{obs}}^{\text{}}$. The proposed IPW-adjusted loss accounts for the distributional shift ($\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$) by up-weighting under-observed points, as described in the following lemma.
@@ 4.2 ICYM2I-learn: Estimating performances for informativeness under missingness
Evaluation. Current works suffer from an analogous bias in model evaluation, by relying on a hold-out set from the observed distribution ($\Omega_{\text{obs}}^{\text{}}$). To estimate a given metric on the true underlying distribution, one must correct this metric using a similar correction as previously described. Li et al. describes how to correct for both AUC and Brier score using IPW.
@@ 4.3 ICYM2I-PID: Partial information decomposition for multimodal informativeness
Under missingness, we have samples from $\Omega_{\text{obs}}^{\text{}}$ instead of $\Omega^{\text{}}$. Estimating PID measures in this setting requires adjusting for the $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$ shift. The three-way mutual information is the key estimand of interest to obtain any PID bound, requiring the following correction (derived in Appendix A):
@@
and analogously for other quantities. Second, these mutual information-based quantities can be equivalently estimated using entropy-based measures (see derivations in Appendix B). Third, to enable PID for high-dimensional data, we parameterize $q$ via unimodal neural networks $f_{1}$ and $f_{2}$, motivated by Liang et al. [46]. Then,
$\Delta_{\Omega^{\text{}}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={p_{\Omega^{\text{}}}}_{\phi}(x_{i},y)\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
in which $\Omega^{\text{}}_{\phi}$ is a re-parametrization of $\Omega^{\text{}}$ using neural networks and learned using samples from $\Omega_{\text{obs}}^{\text{}}$. That is, prior work considers ${p_{\Omega^{\text{}}}}_{\phi}(y,x_{i})\approx f_{\phi}(y\mid x_{i})p_{\Omega_{\text{obs}}^{\text{}}}(x_{i}),\forall i\in\{1,2\}$, with $f_{\phi}$, a unimodal neural network, trained on samples from $\Omega_{\text{obs}}^{\text{}}$ to predict $Y$.
Due to missingness, ${\Omega^{\text{}}}_{\phi}\neq{\Omega_{\text{obs}}^{\text{}}}_{,\phi}$, and should be approximated using IPW corrections for the ${\Omega^{\text{}}}_{\phi}\mapsto{\Omega_{\text{obs}}^{\text{}}}_{,\phi}$. We use the modified Sinkhorn-Knopp algorithm [63] to enforce matching the two-way IPW-corrected marginals to constraint $q\in\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$ [^2] Prior works match conditionals using samples from $\Omega_{\text{obs}}^{\text{}}$ without correction: $q(y\mid x_{i})\approx\Omega_{\text{obs}}^{\text{}}(y\mid x_{i})$ which may produce biased PID bounds.. Using the fact that all PID bounds can be obtained by minimizing the three-way mutual information (see Bertschinger et al. [47], Lemma 4), we summarize our algorithm to minimize $I_{q}(Y:(X_{1},X_{2}))$ in Appendix C.
@@ 5 Experiments
To better understand the connection between performance, information decomposition and missingness, we propose a simulation (detailed in Appendix E), two semi-synthetic studies that reflect real-world missingness mechanisms (see Appendix F), and a real-world case-study.
@@ 5.1 Simulation and semi-synthetic experiments
In Table 2, each point represents the estimated PID value (Unique 1, Unique 2, Shared, and Complementary) for one simulation under the training and evaluation IPW-corrections and the oracle performance, i.e., a model trained and tested on $\Omega^{\text{}}$. Specifically, columns reflect evaluation correction, while rows reflect training correction. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the performance one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. This observation shows that the proposed ICYM2I  best recovers the true informativeness of each modality, despite relying on $\Omega_{\text{obs}}^{\text{}}$. Appendix E echoes the same observation when evaluating model performance.
Observed ($\Omega_{\text{obs}}^{\text{}}$) Underlying ($\Omega^{\text{}}$)
$\Omega^{\text{}}$ [Uncaptioned image] [Uncaptioned image]
The semi-synthetic experiments examine the effect of enforcing increasing missingness on the performance and information decomposition of UR-FUNNY [64] and hateful memes [65], two foundational real-world datasets used in the multimodal literature for affective computing and content moderation. Table 3 summarizes the effect of enforcing 70% missingness on estimating multimodality informativeness across these datasets, demonstrating the generalizability of our proposed strategy across real-world datasets. Appendix F further illustrates the robustness of the methodology under different levels of missingness in these datasets.
Table 3: Impact of 70% missingness on multimodality information for UR-FUNNY [64] and Hateful Memes [65]. Parentheses denote standard deviation across batches.
@@
Text Image/Video Image + Text Unique Text Unique Image Shared Complementary
@@
Structural heart disease (SHD) is a set of conditions that affect the heart’s physiology and is typically diagnosed using transthoracic echocardiograms (TTEs) [66]. However, TTEs are often underutilized in the United States due to diagnostic stewardship and competing financial incentives [67]. Prior work using unimodal models with common modalities in electrocardiograms (ECGs) [68, 69] and chest radiographs (CXRs) [70] has demonstrated that non-TTE modalities can detect structural heart disease labels. However, CXRs are not systematically collected in conjunction with ECGs, leading to systematic missingness patterns. We, therefore, evaluate ICYM2I on this clinical task to evaluate the informativeness of CXRs in diagnosing SHD, despite its missingness.
Dataset. Our study population consists of a retrospective study gathering 98,397 adult patients who received an ECG and a TTE within one year of each other. The population has 20.56% SHD prevalence. In this cohort, 12,587 members (12.79%) have recorded CXRs. For subjects with multiple echocardiograms, we select the first TTE to model opportunistic screening with non-TTE modalities. All data were collected from an academic urban medical system between 2008 and 2022. We generate embeddings using modality-specific foundation models–ECG embeddings are generated using ECG-FM [71] and CXR embeddings with ELIXR-C [72]. Data are split temporally, where subjects with TTEs collected on or after 2018 ($n=40,734$) are allocated to the test set. All data were de-identified, retrospective, and collected for clinical purposes from an academic hospital system, with approval from the Institutional Review Board. Appendix G contains further details regarding preprocessing and embedding generation.
Results. Table 4 presents the performance of each uni- and multimodal model, along with the associated information decomposition. While both the observed and corrected analyses demonstrate the importance of ECG in modeling SHD, the corrected results raise questions about the information gain associated with CXR. Naive decomposition suggests the unique information in CXRs at about 5% of the total information. However, ICYM2I  reduces this unique contribution to 1.8% while increasing estimates of shared information between ECG and CXRs for SHD detection. In contrast to domain knowledge, where ECGs capture electrophysiology while CXRs capture structure and anatomy, two distinct aspects of cardiac health, the corrected complementary and shared results, and low unique information of CXRs suggest that CXRs are not independently useful for SHD diagnosis. Note that our results indicate that the multimodal model performs slightly worse than the unimodal ECG model, reflecting the potential overfitting risk associated with a large number of features.
@@
ECG CXR ECG + CXR Unique ECG Unique CXR Shared Complementary
Observed 0.83 (0.01) 0.72 (0.02) 0.82 (0.01) 0.11 (0.00) 0.01 (0.00) 0.10 (0.00) -0.00 (0.00)
@@ 6 Discussion
This work formalizes the issue of missingness when considering multiple modalities. We emphasize that existing works commonly overlook modality missingness by discarding samples with missingness at the curation stage, or implicitly assume that the missingness mechanism remains stable when a model is deployed in the target environment. Our work formalizes this problem and demonstrates its ubiquity in the multimodality literature. Most critically, prior work ignores the fact that any perceived informativeness of a modality may result in increased rates of data collection, inducing different missingness patterns at deployment. Our work, therefore, introduces ICYM2I, a correction to estimate the information gain associated with a partially observed modality. Our results demonstrate the methodology’s capacity to correct for biases introduced by missingness across synthetic, semi-synthetic, and real-world multimodal datasets. Finally, we evidence the practical utility of this methodology in a healthcare dataset, demonstrating the diverging conclusions that one would reach if ignoring missingness. Our work highlights the critical importance of missingness in multimodal research and urges practitioners to pay particular attention to this issue by systematically collecting data with incomplete modalities and carefully modeling and accounting for missingness to enhance robustness.
@@
Limitations. The key assumption in our work is that modalities are MAR. Practitioners should ensure that this assumption is appropriate for their data. Importantly, MAR is less restrictive than the implicit MCAR assumption made in the multimodal literature. Furthermore, accounting for MNAR requires distributional assumptions that are rarely met or even testable in real-world settings. Additionally, our work is based on Partial Information Decomposition (PID [47]), which focuses on two input modalities. While theoretical advances to extend to multiple modalities remain an open question [73, 74], in practice, practitioners should consider a one-vs-all approach to inform modality informativeness using our method. Like in prior work on PID-based measures of information gain on high-dimensional data [75, 46], the quality of the representations used may impact the measures returned by ICYM2I. We ensure that our probabilistic estimates are calibrated to mitigate any such challenges.

### v3

@@ 4.1 A motivating example
4.2 ICYM2I-learn: Estimating predictive performance under missingness
Training. Under the MAR assumption, i.e., the missingness is fully explained by observed covariates $C$; that is, the probability of a data point being missing depends only on $C$, we propose to train the model with a weighted loss using samples from $\Omega_{\text{obs}}^{\text{}}$. The proposed IPW-adjusted loss accounts for the distributional shift ($\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$) by up-weighting under-observed samples, as described in the following lemma.
@@ 4.2 ICYM2I-learn: Estimating predictive performance under missingness
Evaluation. Existing work suffers from an analogous bias in model evaluation, by relying on a hold-out set from the observed distribution ($\Omega_{\text{obs}}^{\text{}}$). To estimate a given metric under the true underlying distribution, one must correct this metric using a similar correction as previously described. For instance, Li et al. describes how to correct for both AUC and Brier score using IPW.
@@ 4.3 ICYM2I-PID: Partial information decomposition for multimodal informativeness
Obtaining the partial information decomposition bounds corresponds to estimating the distribution $q\in\Delta_{\Omega^{\text{}}}$ that minimizes three-way mutual information (see Bertschinger et al. (2014), Lemma 4), where $\Delta_{\Omega^{\text{}}}$ is the set of all distributions over $(X_{1},X_{2},Y)$ such that the two-way joints between $X_{i}$ and $Y$ match the true data-generating distribution. Formally, one can focus on solving:
$\min_{q\in{\Delta_{\Omega^{\text{}}}}}I_{q}(Y:(X_{1},X_{2}))$
To operationalize this minimization, Liang et al. (2024a) proposes to minimize this quantity using the observed samples:
$\Delta_{\Omega^{\text{}}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(x_{i},y)\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
where $p_{{\Omega_{\text{obs}}^{\text{}}\phi}}$ is a re-parametrization of $\Omega_{\text{obs}}^{\text{}}$ using neural networks. To enforce the marginals of $q$ to match $p_{\Omega_{\text{obs}}^{\text{}}}$, Liang et al. (2024a) proposes an iterative process using a Sinkhorn–Knopp procedure (Knight, 2008).
Under missingness, we observe samples from $\Omega_{\text{obs}}^{\text{}}$ rather than $\Omega^{\text{}}$. To unbias the PID measures in this setting requires adjusting for the $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$ shift. Effectively, our approach corrects the previous optimization to account for the distribution shift induced by MAR missingness.
First, we introduce a corrected mutual information computation to ensure that we optimize an unbiased estimate of the three-way mutual information under $\Omega^{\text{}}$ using samples from $\Omega_{\text{obs}}^{\text{}}$ (see Appendix A for the complete proof).
@@
$I^{\mathrm{IPW}}_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{1-p(m_{1},m_{2})}{1-p(m_{1},m_{2}|x_{1},x_{2},y)}\log{}\left(\frac{p_{\Omega^{\text{}}}(x_{1},x_{2},y)}{p_{\Omega^{\text{}}}(x_{1},x_{2})p_{\Omega^{\text{}}}(y)}\right)\right]$
Then, ICYM2I-PID uses the following projection set to operationalize PID-bound estimation while accounting for missingness:
$\Delta_{\Omega^{\text{}}}$ $^{\textsc{ICYM${}^{2}$I}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={p_{\Omega^{\text{}}}}_{\phi}(y,x_{i})\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
$=$ $\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)=\text{IPW}_{{p_{\Omega^{\text{}}}}_{\phi}}({p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(y,x_{i}))\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
where $\text{IPW}_{q}(p)$ is an IPW correction for the shift $p\mapsto q$ using samples from $p$ and $p_{\phi}$ is a re-parametrization of $\Omega^{\text{}}$ using neural networks and learned using samples from $\Omega_{\text{obs}}^{\text{}}$ via the weighted loss introduced in Lemma 1. Importantly, the proposed correction is agnostic to the parametrisation of $q$.
Estimating PID under missingness therefore consists in optimizing:
$\min_{q\in{\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}}}I^{\mathrm{IPW}}_{q}(Y:(X_{1},X_{2}))$
Similarly to Liang et al. (2024a), the unbiased PID estimation framework consists of the following steps (detailed in Appendix C):
  1. 1.
    Model the missingness mechanism. Train a model to estimate the probability of missingness given $C$ to obtain IPW weights for correcting the distribution shift.
  2. 2.
    Train corrected unimodal and multimodal models. Train each model ($f_{1}(x_{1}),f_{2}(x_{2}),$ and $f(x_{1},x_{2})$) using the IPW-corrected loss introduced in Lemma 1. These models can be trained with flexible inductive biases tailored to the application at hand.
  3. 3.
    Solve PID optimization. Estimate $q\in\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$ minimizing $I_{q}(Y:(X_{1},X_{2}))$, where $q$ is parameterized by the product of two unimodal networks. We note that these mutual information-based quantities can be equivalently estimated using entropy-based measures (see derivations in Appendix B). To match the marginal of $q$ to the estimated probabilities of the corrected unimodal neural networks, we apply a modified Sinkhorn–Knopp procedure (Knight, 2008) using IPW-corrected unimodal distributions [^2] Prior work typically matches $q(y\mid x_{i})$ to $\Omega_{\text{obs}}^{\text{}}(y\mid x_{i})$, which yields biased PID estimates under missingness.. Our framework is agnostic to the choice of parametrization of $q$, provided that calibrated probabilistic scores are learned.
  4. 4.
    Estimate the PID components. Given $q$, compute PID quantities using the bounds of Bertschinger et al. (2014), corrected via the IPW-correction introduced in Appendix A.
@@ 5 Experiments
To better understand the connection between performance, information decomposition, and missingness, we propose a simulation (detailed in Appendix E), two semi-synthetic studies that reflect real-world missingness mechanisms (see Appendix F), and a real-world case study.
@@ 5.1 Simulation and semi-synthetic experiments
Table 2: Comparison between estimated AUC performance under the different training and evaluation corrections and oracle performance on $\Omega^{\text{}}$. $\epsilon$ denotes the RMSE between estimated and oracle AUCs.
Evaluation
Observed Underlying
$\Omega_{\text{obs}}^{\text{}}$ $\Omega^{\text{}}$
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
In Table 2, each point represents the estimated performance value for one simulation under the training and evaluation IPW-corrections and the oracle performance, i.e., a model trained and tested on $\Omega^{\text{}}$. Specifically, columns reflect evaluation correction, while rows reflect training correction. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the performance one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. This observation shows that the proposed ICYM2I  best estimates the performance associated with each modality, despite relying on $\Omega_{\text{obs}}^{\text{}}$. Appendix E echoes the same observation when estimating PID.
The semi-synthetic experiments examine the effect of enforcing increasing missingness on the performance and information decomposition of UR-FUNNY (Hasan et al., 2019) and hateful memes (Kiela et al., 2020), two foundational real-world datasets used in the multimodal literature for affective computing and content moderation. Table 3 summarizes the effect of enforcing 70% missingness on estimating multimodality informativeness across these datasets, demonstrating the generalizability of our proposed strategy across real-world datasets. Appendix F further illustrates the robustness of the methodology under different levels of missingness in these datasets and explores MNAR patterns.
Table 3: Impact of 70% missingness on multimodality information for UR-FUNNY (Hasan et al., 2019) and Hateful Memes (Kiela et al., 2020). Parentheses denote standard deviation across batches.
@@
Text Image/Video Image + Text $_{\text{text}} Unique$$_{\text{image}}$ Shared Complementary
@@
While our core contribution is methodological, this section illustrates how ignoring missingness can lead to biased estimates of the informativeness of a given modality in a real-world setting where modalities are commonly missing. Specifically, we study structural heart disease (SHD), a set of conditions that affect the heart’s physiology, which is typically diagnosed using transthoracic echocardiograms (TTEs) (Writing Committee Members et al., 2021). However, TTEs are often underutilized in the United States due to diagnostic stewardship and competing financial incentives (Papolos et al., 2016). Prior work using unimodal models with common modalities in electrocardiograms (ECGs) (Elias et al., 2022; Ulloa-Cerna et al., 2022) and chest radiographs (CXRs) (Bhave et al., 2024) has demonstrated that non-TTE modalities can detect structural heart disease labels. However, CXRs are not systematically collected in conjunction with ECGs, leading to systematic missingness patterns. We therefore evaluate ICYM2I on this clinical task to assess the informativeness of CXRs for diagnosing SHD, despite their missingness.
Dataset. Our study population comprises a retrospective cohort of 98,397 adult patients who received an ECG and a TTE within 1 year of each other. The population has 20.56% SHD prevalence. In this cohort, 12,587 patients (12.79%) have recorded CXRs. For subjects with multiple echocardiograms, we select the first TTE to model opportunistic screening with non-TTE modalities. All data were collected from Columbia University Medical Center, a large academic medical system in New York City, between 2008 and 2022. Data are split temporally, with subjects with TTEs collected on or after 2018 ($n=40,734$) allocated to the test set. All data were de-identified, retrospective, and collected for clinical purposes within an academic hospital system, with Institutional Review Board approval. Appendix G contains further details regarding preprocessing, embedding generation, and the ICYM2Iimplementation.
Results. Table 4 presents the performance of each uni- and multimodal model, along with the associated information decomposition. While both the observed and corrected analyses demonstrate the importance of ECG in modeling SHD, the corrected results raise questions about the information gain associated with CXR. Naive decomposition suggests the unique information in CXRs accounts for about 5% of the total information. However, ICYM2I  reduces this unique contribution to 1.8% while increasing estimates of shared information between ECG and CXRs for SHD detection. In contrast to domain knowledge, where ECGs capture electrophysiology while CXRs capture structure and anatomy, two distinct aspects of cardiac health, the corrected complementary and shared results, and low unique information of CXRs suggest that CXRs are not independently useful for SHD diagnosis. Note that our results indicate that the multimodal model performs slightly worse than the unimodal ECG model, reflecting the potential risk of overfitting associated with a large number of features.
@@
ECG CXR ECG + CXR $_{\text{ECG}} Unique$$_{\text{CXR}}$ Shared Complementary
Observed 0.83 (0.01) 0.72 (0.02) 0.82 (0.01) 0.11 (0.00) 0.01 (0.00) 0.10 (0.00) 0.00 (0.00)
@@ 6 Discussion
This work formalizes the issue of missing modalities in multimodal settings. We emphasize that existing work commonly overlooks missingness by discarding samples with any missing modality during curation, or implicitly assumes that the missingness mechanism remains stable when a model is deployed in the target environment. Our work formalizes this problem and demonstrates its ubiquity in the multimodality literature. Most critically, prior work ignores that any perceived informativeness of a modality may result in increased rates of data collection, inducing different missingness patterns at deployment. Our work, therefore, introduces ICYM2I, a correction to estimate the information gain associated with a partially observed modality. Our results demonstrate the methodology’s capacity to correct for biases introduced by missingness across synthetic, semi-synthetic, and real-world multimodal datasets. Finally, we demonstrate the practical utility of this methodology on a healthcare dataset, showing the divergent conclusions that one would reach if missingness were ignored. Our work highlights the critical importance of missingness in multimodal research and urges practitioners to pay particular attention to this issue by systematically collecting data with incomplete modalities and carefully modeling and accounting for missingness to enhance robustness.
@@
Limitations. The key assumption in our work is that missingness is MAR. No theoretical guarantees exist under MNAR patterns. While distinguishing these assumptions is empirically untestable, practitioners should ensure that this assumption is appropriate for their data. Importantly, MAR is less restrictive than the implicit MCAR assumption made in the multimodal literature, and does not require distributional assumptions that one must assume to tackle MNAR patterns. Additionally, our work is based on Partial Information Decomposition (PID) (Bertschinger et al., 2014), which focuses on two input modalities. In practice, practitioners could consider a one-vs-all approach to inform modality informativeness using our method. However, extending the decomposition to more than two modalities remains an open challenge (Griffith and Koch, 2014; Kolchinsky, 2022), where notions of mutual information are not well-defined beyond three-way mutual information. As in prior work on PID-based measures of information gain on high-dimensional data (Liang et al., 2023; 2024a), the quality of the representations used may impact the measures returned by ICYM2I. We ensure that our probabilistic estimates are calibrated to mitigate such challenges. Additionally, PID is designed for datasets with paired modalities, where the input modalities are associated with an instance (e.g., a patient and their associated chest X-ray and electrocardiogram); consequently, our method does not extend to unpaired settings. Finally, our proposed method relies on a notion of instance for which all modalities are observable. Extending this method when there is no notion of an instance, i.e., unaligned modalities, could be considered but would require different inductive biases to model the underlying unimodal and multimodal probabilities.
Ethics statement. Our work demonstrates the impact of missingness on performance estimates in multimodal learning. We demonstrate the utility of our method in a crucial healthcare use case. However, the methodology remains a proof of concept that would require additional testing to be deployed in a real-world context. Our study is approved by the Columbia University Medical Center Institutional Review Board (IRB-AAAU7973). We have extracted all data in HIPAA-compliant servers, and our experiments are also conducted on HIPAA-compliant compute despite being deidentified for extra caution. While beyond the scope of this work, modality completeness is not uniform across demographic subgroups and can manifest in data collection policies, such as differential access to care based on insurance status. Our method could provide important insights into the utility of multimodal predictions in such settings.
Reproducibility statement. Theoretical proofs are provided in Appendix A. All code for applying the proposed ICYM2I  and reproducing all synthetic and semi-synthetic results presented in this work is publicly available on Github333 https://github.com/reAIM-Lab/ICYM2I. A summary of the computational resources required to reproduce our results is given in Appendix H.